OpenStudy (anonymous):

Are the lines y = –x + 9 and 5x – 5y = 10 perpendicular? Explain.

OpenStudy (dan815):

yes because if you were to divide this equation by 5 and rearrange in for y, you can see that the slope is the negative reciprocal of the slope of the first 5x – 5y = 10 x – y = 2 y=x-2, slope=1 which is negative reciprocal of -1

OpenStudy (dan815):

The slope must be a negative reciprocal, you can see from this diagram below

OpenStudy (dan815):